OpenStudy (anonymous):

will someone please help me with these 9 problems that i have already did and got wrong... Please help

OpenStudy (anonymous):

these two i had to find the base area... and i got them wrong

OpenStudy (anonymous):

What are the shapes of bases on each picture? What are the formulas to calculate areas of these figures?

OpenStudy (anonymous):

traingle and a circle

OpenStudy (anonymous):

Correct. What kind of triangle? It's a *regular* pyramid.

OpenStudy (anonymous):

equallaterial

OpenStudy (anonymous):

Yup. So you have to find an area of a circle given it's radius and an area of an equilateral triangle given the length of this side. Both formulas can be found easily, e.g. on wikipedia. If you want an additional challenge you can devise the formula for triangle using the Pythagorean theorem (if you draw its height you'll get a right triangle).

OpenStudy (anonymous):

for the triangle wouldnt it be 1/2*4*4=8cm

OpenStudy (anonymous):

Nope, its height does not equal 4.

OpenStudy (anonymous):

what is the height

OpenStudy (anonymous):

"Altitude" is another word. http://en.wikipedia.org/wiki/Altitude_%28triangle%29 I assume you used the formula for triangle that's given as area = 1/2 * base * height.

OpenStudy (anonymous):

yea... for the cone the base area would be 113.0976 or 113.1in

OpenStudy (anonymous):

\[113.1 in^{2} \] That is square inches. Correct.

OpenStudy (anonymous):

i still dont know how to find the base area of the pyramid

OpenStudy (anonymous):

Area of base: ½b×h

OpenStudy (anonymous):

1/24*7=14cm

OpenStudy (anonymous):

1/2*4*7=14cm

OpenStudy (anonymous):

The wikipedia page about equilateral triangles has the formula that uses only the side length.

OpenStudy (anonymous):

1/2*4*7=14cm so this isnt right

OpenStudy (anonymous):

No, it's not right.

OpenStudy (anonymous):

i give up

OpenStudy (anonymous):

Too early. You're given an equilateral triangle with side length of 4cm. Apply the formula from: http://en.wikipedia.org/wiki/Equilateral_triangle

OpenStudy (anonymous):

A=\frac{\sqrt{3}}{4} a^2

OpenStudy (anonymous):

Yes, this one.

OpenStudy (anonymous):

6.93

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

okay.. how are these two wrong.. finding the volume

OpenStudy (anonymous):

In both cases you got the base area wrong. First case is the same as the problem you've just solved. Second one has a base that can be thought of as consisting of multiple triangles.

OpenStudy (anonymous):

the first one the base area would be 10.83

OpenStudy (anonymous):

Indeed.

OpenStudy (anonymous):

so the volume would be 32.49in cubed

OpenStudy (anonymous):

True.

OpenStudy (anonymous):

what formula would i use to get the base area for the hexagon

OpenStudy (anonymous):

It consists of six equilateral triangles.

OpenStudy (anonymous):

\[\sqrt{6}\div4\times3^{3}\]

OpenStudy (anonymous):

\[3^{3}?\]

OpenStudy (anonymous):

7^3

OpenStudy (anonymous):

Let's do it more slowly. What's the area of one of the triangles that form that hexagon?

OpenStudy (anonymous):

\[\sqrt{3}\div4\times7^{3}\]

OpenStudy (anonymous):

7^2

OpenStudy (anonymous):

Yes. And there's six of them. So the area of the whole hexagon is..?

OpenStudy (anonymous):

21.21*5

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

21.22*6

OpenStudy (anonymous):

Much better. You can get the correct volume now.

OpenStudy (anonymous):

127.31

OpenStudy (anonymous):

Looks good to me.

OpenStudy (anonymous):

381.92 would be the volume?

OpenStudy (anonymous):

No, you must have missed something in the formula for volume.

OpenStudy (anonymous):

127.31 this is the area of one side base or all

OpenStudy (anonymous):

There's only one base in this case.

OpenStudy (anonymous):

127.31 is the base area right

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

than to get the volume dont i * by 3

OpenStudy (anonymous):

You do, but there's a division there too.

OpenStudy (anonymous):

190.97

OpenStudy (anonymous):

Pyramid (1/3) * (base area) * height

OpenStudy (anonymous):

That's the correct formula, yes. Height is 3cm if I read the picture correctly.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

(1/3) * 3 * 127.31 = ..?

OpenStudy (anonymous):

190.97cm^3

OpenStudy (anonymous):

127.32 * 3 / 3 = ..?

OpenStudy (anonymous):

127.32

OpenStudy (anonymous):

Aye.

OpenStudy (anonymous):

(127.32 * 3) / 3 = 127.32

OpenStudy (anonymous):

Multiply by the numerator, divide by denominator. That's how you multiply a number by fraction. Surprisingly 127.32 cubic centimeters is the answer. It's the same number as the base area, but unit differs (cubed vs squared).